Find all forces in Members AB AH BH BC CD CF GF(a) Determine the forces in members: AB, AH, BH and BC using Method of Joints. (b) Determine the forces in members: CD, CF, a
Answer #1

G 4.5m H 3m IB ARBE in LORN PRIN GKN 3m 3 m 3m 3m ZFy=o- RA TRE - 6-8-2=0 RA TRE = 16 EME=0 RAX (12) - 649 - 846 - 2a3=0 RA=using method of Joint at at Joint A Fan é 3/45 tano = 3 3 o=tan(1) -45 FAB RA=9kn & Fy=0 FAH sino +RA = 0 Fan sin 45 +9=0 Fat Joint B. и FAB FBC B 6 KN & Fy=0 Fви — бе о Fви – GEN (tension) & fx=0 с – ве = (tension) [Faca Fire - 9 KN ам Au Хиш АВ |X 4.5m H 3m D A7KN=RE B LUN 6 KN 8W 9kn 3m 3m 3m using method of section tano 3 4.5-3 453 FGF &= tan (1) A FLE 3m B ECD $ =& Fx=0 For coso t FCF cos o + FLD =0 ForF cos 26.565 t FCF cos 45 +7=0 FGEVO cos 26.365 t Feu cos 45=-7 -i) E fy=0 FGF sin 2plz comment if you have any doubts. Plzz upvote if you like the solution